1st half of night: NON-sidereal SW testing: 10 December 2011 UT

Observer: JHill (LBTO), CWoodward, KBoutsia (Tucson)
Telescope Operator: SAllanson (LBTO)
Software Support: MDelapena, CBiddick, NCushing (Tucson)
Telescope Support: JUrban (LBTO)
Instrument Support: TShih (LBTO)

2nd half of night: LBC focal plane flatness test; m2 astigmatism correction data and test, adjustments for SX AO m2 for collimation model, adjustments for LBBG pointing model: 10 December 2011 UT

Observer: JHill (LBTO), Arakich, DThompson (Tucson)
Telescope Operator: SAllanson (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: TShih (LBTO)

Summary

Light cirrus at the beginning of the night.

Non-sidereal observing had some success, but also a fair number of problems. Guiding for non-sidereal LBC targets works, but guiding for MODS does not work.

Andrew got various tests and data.

We got a 3 hour headstart in the LBBG work for tomorrow.

Preparations

Details

00:15 5Bino_Bias

00:29 Slew for flats B:g-sloan - R:r-sloan
  • 1sec is saturated
  • try 0.5sec

00:43 start flats g-r : 0.8sec PA=0
  • 1.5sec - PA=180

00:54 Slew to WT10_278 for collimation and rangebal
 2381    149   -865  -2216    -61    124     52   1.74   1.40     red
 -1848     47   -464  -1202   -520    567     57   1.45   1.01    blue

 -1114    -52    293    647     28    208     68   1.36   0.56     red
-262   -126    138    -82     25    122     74   2.01   0.17    blue

-1349    -44     47   -369   -419    208     52   1.79   0.61     red
-204    102   -240   -324   -155     87     55   2.35   0.22    blue

603     61    132    -52    372     18     21   1.27   0.31     red
-294   -295    -91    229     89    121     10   1.91   0.21    blue

 113    -72     66     56      5   -102      0   1.18   0.12     red
 535   -193     44     19    -54   -103     -9   2.22   0.24    blue
  • dimm 1.2''

01:06 Slew to the focus field for CUNO
  • focus-cuno
  • dofpia and rangebal again
  357    108    -41    363    219   -120      8   1.16   0.26     red
 -10    -30    144    -45   -136     87     13   1.51   0.12    blue

01:15 first we create the ephemeris with the NSIGUI
  • IIF NON-sidereal ON
  • send CunoLBC.ob from the LBC interface
  • the guide rates for the LBC were 90% of what they are supposed to be... do we need to fix that?? - talk with JHill
  • the weight factor should probably be 1.0 for non-sidereal observing (different for fast and slow moving sources?)

01:27 repeat procedure to repeat OB
  • IIF NON-sidereal ON
  • send CunoLBC.ob from the LBC interface
  • no guide rate changes have been applied yet...
  • we don't see the object in the field

01:37 repeat procedure with 60sec of integration to see if the stars in the image are creating trails.
  • they do seem elongated - also the target is slightly elongated (less than the rest though)
  • we do lbckill/lbcstart

01:46 changed guide rate factor to 0.98 to decrease the ellipticity on the asteroid that
  • it looks slightly elongated again but a bit better than before
  • the target is saturated (60sec integr) and we cannot really tell if it is actually rounder than the field stars
  • we need to go to a fainter target

01:52 2003sb45_dec10
  • NSIGUI name of the ephemeris file is : 143658_2011-Dec-10.txt

02:03 dofpia
-1418     27    157    -26      7   -279    -61   0.77   0.59     red
-501    113   -115    173    419   -267    -74   1.11   0.34    blue

  -91    -49    185    339     32    -32    -11   1.05   0.19     red
 -354     -7    -90   -101     -2     94     -7   1.48   0.18    blue

02:13 SLew to a new target (C/2010 G - Hill)
  • too close to the moon - change targets
  • NSIGUI file C2010G2_2011-Dec-10.txt

02:20 Slew to a target away from the moon (C/2010 S1 (LINEAR))
  • dofpia
 -218     51     68    214     62   -134    -57   0.70   0.18     red
  326   -143    -57    568   -114   -248    -57   1.03   0.34    blue

 -391     18     64    -50      9     19    -19   0.71   0.18     red
 -214     79     -9     35    -15    -14    -20   1.05   0.12    blue

  101     75     15     56      7     21     -9   0.76   0.10     red
 -301     43    102    -65    -11    115      8   1.19   0.17    blue
  • no way to half the dofpia exptime
  • since we are pointing into the galactic plane and there are too many pupils and there is no good measurement for sky background

02:30 NON-sidereal procedure for 2010-S1
  • NSIGUI file C2010S1_2011-Dec-10.txt

02:40 We added the z-filter because we were too close to saturation
  • we repeat the procedure for 2010-S1
  • it seems to be working - when subtracting 2 blue images we see all the stars move and the comet disappears....
  • also in the last red image with the z-filter all stars have obvious trails

02:56 Slew to a fainter target - C/2008 FK75 (L-SS)
  • dofpia
 298      8    154    -58    136   -162    -15   0.95   0.19     red
 628   -177    151     39   -189    -90    -15   1.16   0.29    blue

 -110    -93     74     32    -33    -12      4   0.88   0.10     red
 82   -195     64    -21    -51    -46     12   1.25   0.11    blue
  • press set in the NSIGUI and then send the OB
  • 2008-FK.ob
  • in the image display the source seems saturated but when we run imexam the counts go up to 8000counts - this is because it is background subtracted...

03:18 repeat OB - 2008-FK.ob just to be sure that we are not saturating...

03:26 dofpia and then we repeat 2008-FK.ob with the standard 4-point dither
 -802    -74    -13     17     78    -64    -18   0.80   0.33     red
-292   -109    105     76    125     26     -8   1.20   0.16    blue

  42     66    130    -79     71    -14     -9   0.95   0.10     red
 -216     17     73    -86    -24    122     11   1.24   0.15    blue
  • press set in the NSIGUI and then send the OB
  • 2008-FK.ob - with dither

03:38 Abort because it seems not to have the non-sidereal tracking flag after the first offset position...
  • probably not set the NONsidereal flag before sending this preset

03:40 repeat OB ...
  • press set in the NSIGUI and then send the OB
  • 2008-FK.ob - with dither

03:51 LBC error - blue camera set to STOP on the power control panel (7min ago)
  • OB aborted
  • just turning ON and OFF the camera, does not work -- "command rejected because another one is running."
  • lbckill - lbcstart - 8min lost

03:59 since we are still in non-sidereal mode we need to send the preset again and then dofpia
-1647     89     35     47   -162   -113      9   0.83   0.66     red
 -924    -90     43    -47     80    -75    -14   1.09   0.38    blue

 -56     90     36     62    -93     30    -21   0.82   0.10     red
 433     23    132     77    115    -93    -24   1.16   0.21    blue

  245    105     19    -15     79     -8     -5   0.83   0.13     red
 35    160    -44    -29    -40    -51    -20   1.07   0.10    blue

04:04 NSIGUI - set!
  • 2008-FK.ob
  • 15min lost if you consider that we had to focus again - re-acquire the target and repeat the first exposures

04:14 abort again because something is wrong with the PCS * the first image after the dither is not guiding - we have long trails * the second image after the offset for the blue side seems to be hopping (the same star in 3 different positions - no continuous trail - 2 jumps) * we hit a limit on the left side and then something else strange happened... this thing needs further investigation and we decide to switch to MODS.

04:17 SWITCH TO MODS

04:58 first MODS target
  • Set NSIGUI - the txt file is 1983TB_2011-Dec-10.txt
  • acqMODS 3200P.acq
  • it gave an error about the guide probe not being able to reach the position - the guide star was wrong
  • we will also need to unwrap very soon - so we do this before sending the preset again

05:06 send the acquisition again
  • set NSIGUI - acqMODS 3200P.acq
  • the GS has been updated
  • it gave a warning that guiding was not stable after 6 iterations - the star on the guider is non the upper left corner

05:14 acqMODS 3200P.acq
  • we do this in sidereal mode just to collimate because before it was sending wrong corrections
  • the SlitGO lines have been commented

05:19 send the NON-sidereal acquisition - preset TRACK
  • acqMODS 3200PnonSide.acq mods1r.20111210.0002.fits - *.0003.fits
  • cirrus is passing by...
  • we cannot run modsAlign in remote - JHill aligned the mask for us from the mountain

05:41 the source is not in the slit - probably because a long time has passed to trace the problem
  • we repeat acquisition mods1r.20111210.0006.fits - *.0008.fits
  • we exited and entered LBTO with ssh -X but it did not work - we got the same error
  • we ask JHill again to align the slit

05:51 the source is centered in the slit and we start the integration
  • execMODS 3200P.obs mods1r.20111210.0009.fits - *.0011.fits / mods1b.20111210.0001.fits - *.0003.fits
  • we take the slit and prism out and we take a 10sec image to see how much the source has moved mods1r.*.0012.fits

06:07 go to a slower moving target C/2010-S1
  • much more cirrus than before
  • since now it is not moving very fast, we should not have to make an active preset with sidereal guiding and then a track preset with Non-Sidereal
  • we just send the normal acq script with non-sidereal guiding - the guide star is drifting outside the guider
  • we abort and we send the preset again

06:13 acqMODS c2010s1linearNonSide.acq

06:19 send a normal sidereal active preset
  • acqMODS c2010s1linear.acq
  • failed because we did not have the correct GS in this script
  • update GS and repeat acquisition - it still cannot guide
  • preset failed because of instrument error
** ERROR: Command 'INSTCONFIG red imaging' timed out after 120 seconds
** Abort, Retry or Ignore? > Retry

** c2010s1linear.acq aborted with errors at line 13 command 'INSTCONFIG red imaging'
  • this time is an instrument error.

06:29 send the preset again changing the GS with a brighter one
  • anyway, it seems that the guider is not working properly
  • the star in the guider looks bad anyway...
  • probably best to move to something higher

06:32 the script gave error again
** ERROR: INSTCONFIG instrument configuration already in progress - command disallowed
** Abort, Retry or Ignore? > Abort

** c2010s1linear.acq aborted with errors at line 13 command 'INSTCONFIG red imaging'
  • the CCD readout box seemed to be stuck - we manually changed readout modes and now it seems to be set properly

06:41 in order to create the 60P NSIGUI input file we write 60P in the Target Name
  • it returns error - No such object
  • we look at the log - "View Log" and if it has multiple entries then you choose the correct one for your target
  • for 60P in 2011 the number for the NSIGUI is 900605 and the name of the file with the ephemeris is 900605_2011-Dec-10.txt

06:44 acqMODS 60p.acq - normal sidereal acquisition in order to collimate
  • preset failed
>> PRESET ACTIVE
** ERROR: PRESET IIF error - PresetTelescope result status: Error, Stop[122] complete, PCS SetEquatorialTarget co-pointingLimitError performed. Target on side left is greater than 9.44121PCS SetEquatorialTarget failed on side left., PCS SetEquatorialTarget failed on side left.
  • steve is sending the clear preset from the command line

06:49 send the preset again in sidereal mode
  • no star in the guider - clouds???
  • preset aborted

06:55 send the preset again in NON-sidereal mode (preset active)
  • we do not see the star in the guider
  • MODS though, thinks that everything is fine and it goes on taking images
  • close hatch in the middle of an integration (07:00)

07:01 END OF NON SIDEREAL TESTING

ARakich, DThompson

LBC UI found unresponsive (all clicks return "command rejected since another is running). Did a kill/start on lbccontrol.

7:27 Preset to WT10_214 to focus/co-point

7:35 converged in a few minutes. Rangebalanced.

7:45 Running Olga's OB to take extra and Intra-focal pupils at

7:46 Running @ PA= 0 (RB_rVinex4.ob)

7:49 Running RB_rVexin4_PA180.ob

7:57 redid FPIA, running RB_rVexin4.ob folllowed by RB_rVinex4_PA180.ob

8:01 DOFPIA

8:06 Running RB_rVexin4_PA180.ob followed by RB_rVinex4.ob

8:14 DOFPIA

8:18 converged now doing superfoc

8:19 running RB_rV_superfoc

8:23 running RB_rV_superfocPA180

8:30 DOFPIA

8:35 running RB_rV_superfocPA180

8:39 running RB_rV_superfoc

8:43 DOFPIA

8:47 running RB_rV_superfoc

8:51 running RB_rV_superfocPA180

8:55 Reconfiguring for MODS1

Verified MODS guiding with the 50 pixel guide box, and without the bright edge.

WT10_269

09:15 Adjusted X-Y position of M2 to eliminate coma. Xglobal=-1.2 Yglobal=0.5 Zglobal=-1.000 at T=-2 degC
Also needed -6000 nm Z6.

09:43 delta IE +7.8 delta CA +3

Manual Astigmatism Measurements

Taking Calibration images for Quick Astigmatism Correction

Taking some images with known steps in Z4 and Z6 on GS 8 WT10_269 for analysis early next week. After these images are taken, handing over to John to begin LBBG focal plane comissioning work.
1) Collimate on GS 8.

2) Now, after stopping collimation and guiding, I want images taken with
the each of the following aberrations loaded into M1 PSF Gui: Send whatever
preset you work out is best to take these images:

3) Send to PSF M1 GUI + 20000 Z4

4) Take images with Z6 added in steps of +2000 nm from 0 nm to 20000 nm.

5) Recollimate

6) repeat 3) - 4) but this time with +10000 nm Z4 added initially.

Images are in /OldRepository/AGW_Data/20111210_MODS1/

Collimating for a while off-axis on gs=8

10:40 ACQUIRE preset on-axis to WT10_269

10:42 Send +20000 Z4 to SX M1, on-axis star still too bright. So return to gs=8 for images

10:44 stopAcquire left

10:45 readGuideCam left -e 2 (obviously too faint)

First image is 0003 (ignore 1 and 2) - note the "interesting" vignetting of this pupil

10:48 Z6=+2000

guider 0004

10:49 Z6=+4000

guider 0005

10:50 Z6=+6000

guider 0006

.......... steps of 2000

Z6=+20000

guider 0013

10:53 recollimate to remove the aberrations

10:55 ACQUIRE preset

Z4=+10000

10:58 stopAcquire left

guider 0014

Z6=+2000

guider 0015

.......... steps of 2000

Z6=+20000 (total)

guider 0024

Check of vignetting in the MODS field

11:04 Apply +20000 Z4 (total)

11:05 ACQUIRE gs=13 but a little faint

11:07 ACQUIRE gs=7

guide 0025

11:11 ACQUIRE gs=5

guide 0026

11:12 ACQUIRE gs=29 (farther out)

guide 0027

Preliminary conclusion is that the vignetting is independent of field position and angle. Must be in the guider path of MODS1 AGW or something like that.

Reconfigure to LBBG

11:15 Reconfigure telescope to start the LBBG work early (How did John let Andrew talk him into this?)

11:25 GCSL hangs while initializing WFS AzCam

Kill/Start GCSL - didn't help

Kill GCSL, stopAGW/startAGW, Start GCSL - didn't help

11:34 Kill GCSL, stop and restart the AzCam windows (without rebooting azcamserver) - WORKED! (same as yesterday)

11:40 ACQUIRE WT10_288
movie image didn't arrive within timeout - aaaaargh

11:43 Stop GCSL, restart GCSL

11:44 ACQUIRE WT10_288 - it worked. (image, but no star)

11:46 ACQUIRE M5_0328 (a wide binary)

11:50 ACQUIRE WT10_288 - centered at IE=-33 CA=-47 at EL=75

11:52 ACTIVE off-axis - seeing has improved to 0.5 arcsec, but image is looking a bit triangular.

12:00 ACTIVE BS9126 gs=0 PA=0 This pupil is at 386,196 (central hole is balance on a particular lenslet) and Doug would like us to move it to 377, 199.

Increase Tip on M3 by 200 arcsec in steps of 20. This doesn't make any obvious motion of the pupil, which means I need a better way to measure it. Blink images wfsc 0105 (before the tip) and wfsc 0120 (after the tip) - this shows that the pupil image moved up perhaps 6 pixels more than needed.

12:22 Decrease Tip to 100 in steps of 20. See wfsc 0131. Absolute Tip is -815 arcsec

12:26 Decreased Tilt by 200 in steps of 20. See wfsc 0137. Absolute Tilt is +680 arcsec. This moved the pupil left toward smaller X (correct direction).

12:31 Decreased Tilt to -400 in steps of 20. See wfsc 0148 for PA=0. Absolute Tilt is +480 arcsec.

12:35 ACTIVE BS9126 gs=0 PA=90 - but there is no star because the astronomer moved M3.

12:39 ACQUIRE M5_0328 IE=-19 CA=-25

12:42 ACTIVE BS9126 gs=0 PA=90 - wfsc 0153

12:44 ACTIVE BS9126 gs=0 PA=180 - (seeing worsens) - wfsc 0156

12:48 ACTIVE BS9126 gs=0 PA=270 - wfsc 0159

The bad news is that there is still pupil wobble when we rotate the rotator based on blinking the images in ds9. The good news is that it looks like we might be within the range of adjustment.

Updated AO_IRTC_L.cfg:
WFScam_hotspot_x	376.7		double	# ccd x coordinate of pixel that represents leftmost x-center of WFS spots 378.0
WFScam_hotspot_y	199.1		double	# ccd y coordinate of pixel that represents leftmost y-center of WFS spots 195.5

13:00 Restart GCSL to get new parameters.

13:02 ACTIVE BS9126 gs=0 PA=0 -

Testing the idl machinery (and my ability to operate it at 6AM)

agw_set_station, 'IRTC'

/OldRepository/AGW_Data/20111210_AGW1/field/

first wfsc image is /Repository/GCSfiles/dailyimages//left_wfscimage000193.fits

Seeing is about 1 arcsec. Sky is mostly clear as the cirrus are leaving just in time for sunrise.

13:15 field_collect (using BS9126_new)

but he forgot iif_register so the presets didn't go anywhere.

Close

13:28 Close for impending dawn and impending sleep.

-- KonstantinaBoutsia - 10 Dec 2011
Topic revision: r11 - 10 Dec 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback